The total earnings of a worker are represented by E = 100 + $10(24−L), where E is earnings and L is the number of hours of leisure. How much will the worker earn if he takes 14 hours of leisure per day? Show solutions.


1
Expert's answer
2022-02-11T10:01:31-0500

Given that, "E = 100 + 10(24\u2212L)," where E is earnings and L is the number of hours of leisure, we need to determine the worker's earnings if the worker takes a leisure of 14 hours per day.

To determine the worker's earnings, we shall substitute for "L=14" in the given equation. So, when "L=14",

"E=(100+10(24-14))=100+100=\$200".

Therefore, when a worker takes a leisure of 14 hours, he earns "\$200".